# BottleHumor: Self-Informed Humor Explanation using the Information Bottleneck Principle

### v1

@@ 1 Introduction
Several datasets for multimodal humor understanding tasks were proposed, where models are tasked with generating free-text humor explanations for an image and a caption Hwang and Shwartz (2023); Hessel et al. (2023); Nandy et al. (2024). However, they are often overlooked in vision-and-language models (VLMs) evaluations, possibly due to the subjective nature of humor and the challenges in evaluating free-text explanations. With that said, VLMs have demonstrated remarkable visual reasoning capabilities on datasets requiring scientific knowledge Lu et al. (2022), commonsense knowledge Schwenk et al. (2022), and spatial reasoning Liu et al. (2023) and there is a prominent line of work on enhancing multimodal reasoning Zhang et al. (2024); Mitra et al. (2024); Mondal et al. (2024); Hu et al. (2024).
@@ 2 Related Work
CoT has been adapted to the vision and language setting Zhang et al. (2024) by adding external knowledge Mondal et al. (2024), extracting a scene graph Mitra et al. (2024), or using visual sketches as intermediate reasoning steps Hu et al. (2024). Most existing works focus on benchmarks such as ScienceQA Lu et al. (2022) and visual commonsense reasoning Schwenk et al. (2022), with (a) definitive/objective answers; and (b) simple evaluation metrics (e.g., ScienceQA is multiple-choice). We focus on multimodal explanation generation tasks in which the answers are open-ended and nuanced. As in CoT, we elicit intermediate reasoning steps from the models, but propose a novel method using the information bottleneck principle to guide generation and selection of useful knowledge for a correct explanation.
@@ 3.1 The Information Bottleneck Principle
where $I$ denotes mutual information, and $\alpha$ is a parameter to balance compression term $I(S,\hat{S})$ with relevance term $I(\hat{S},Y)$.
@@ 3.2 Eliciting Multi-Hop Implications
To guide implication selection for generating the correct output, we provide the image $I$ and caption $C$ to the VLM to generate a set of candidate explanations at each hop: $R^{h}=\{r_{1}^{h},r_{2}^{h},\dots,r_{k}^{h}\}$. One candidate explanation acts as an initial hypothesis, refined iteratively when additional information (implications) becomes available. In the first hop, we generate candidate explanations by providing the VLM with the image $I$, caption $C$, and descriptions $D$. From the second hop onward, we condition—in addition to the previous inputs—on each of the $k$ implications selected in the previous hop (§3.3) to generate $k$ candidate explanations. The prompts used for generating image descriptions, implications, and candidate explanations are in Appendix E.
@@ 3.3 Selecting and Refining Useful Implications
The relevance term is used to ensure that implications provide useful information for generating a target explanation. Since our method is unsupervised, we use the VLM to generate candidate explanations at hop $h-1$: $Y=\{R^{h-1}\}_{i=1}^{I}$, which we use as a proxy for the gold standard answer in the next hop $h$. We measure the relevance of the target implication $P_{j}^{h}$ as the maximum probability (minimum cross entropy loss) for predicting the candidate explanation from the current (textual) inputs $\hat{Z}_{j}^{h}=\{C,D,P^{h-1},P_{j}^{h}\}$, which include the caption, image description, implications from previous hops, and the target implication:
@@
where $\alpha$ is a hyperparameter that controls the trade-off between the compression and relevance terms. In our experiments, we set $\alpha=0.7$.
@@ 3.4 Generating Final Answer
We used Sentence Transformer [^2] BAAI/bge-large-en-v1.5 for all sentence embeddings. The prompts for generating multi-hop implications and explanations are in Appendix E.
@@ 4.3 Baselines
ZS generates a final explanation directly from the image and caption using VLM. CoT follows a similar setup but instructs the model to produce intermediate reasoning chains Wei et al. (2022). Additionally, we implement SR, a multimodal variant of self-refinement Madaan et al. (2023), where a generator produces a response, and a critic evaluates it based on predefined criteria. The critic’s feedback helps refine the output iteratively [^4] Refinement steps set to 2 for fair comparison.. Evaluation criteria include correctness, soundness, completeness, faithfulness, and clarity (details in Appendix G). SR-noC functions identically to SR but without a critic model, refining candidate explanations without feedback. This also serves as an ablation of the implications from BottleHumor. Prompts for baselines are in Appendix G.
@@ 4.4 Evaluation Metrics
For recall, we decompose the reference $ref$ into atomic facts: $\{y_{1},y_{2},...,y_{n}\}$ and check whether each appears in the predicted response $pred$.
@@
where $n$ is the number of atomic facts in $ref$.
Precision follows the same process in reverse, decomposing $pred$ into a list of atomic facts: $\{x_{1},x_{2},...,x_{m}\}$ and verifying their presence in $ref$:
@@
where $m$ is the number of atomic facts in $pred$. Both decomposition and verification use Gemini-Flash-1.5 with a temperature of 0.2.
@@
To assess the reliability of our metrics, we conducted a human evaluation on 130 random samples across all models and datasets via CloudResearch (details in Appendix D). Human annotators determined whether each atomic sentence appeared in the corresponding text (e.g., reference). The average agreement between the LLM-based evaluator and two human annotators was 77.1% ($\kappa=54.1$), similar to the agreement between the two annotators: 75.4% ($\kappa=50.8$), indicating considerable alignment with human judgment. Prompts are in Appendix F.
@@ 5.2 Contribution of Individual Components
Table 2 presents an ablation study where only a single input is provided after refining implications and candidate explanations. GPT-4o and Phi perform better with both inputs, suggesting they effectively integrate relevant information from both to generate improved explanations. In contrast, Flash-1.5 and Qwen2 models rely more on the candidate responses, which contain more readily-useful information than the implications, indicating these models are less proficient at ignoring noisy or irrelevant implications.
@@ 5.3 Assessment of the IB Framework
Table 3: Precision, Recall, and $F_{1}$ scores on intermediate explanations across hops. h stands for hop.
@@ Appendix D Details on human anntations
Appendix E Generation Prompts for Selection and Refinement
@@
Appendix F Evaluation Prompts
@@
Appendix G Prompts for Baselines

### v2

@@ 1 Introduction
Several datasets for multimodal humor understanding tasks were proposed, where models are tasked with generating free-text humor explanations for an image and a caption Hwang and Shwartz (2023); Hessel et al. (2023); Nandy et al. (2024); Hu et al. (2024b). However, they are often overlooked in vision-and-language models (VLMs) evaluations, possibly due to the subjective nature of humor and the challenges in evaluating free-text explanations. With that said, VLMs have demonstrated remarkable visual reasoning capabilities on datasets requiring scientific knowledge Lu et al. (2022), commonsense knowledge Schwenk et al. (2022), and spatial reasoning Liu et al. (2023) and there is a prominent line of work on enhancing multimodal reasoning Zhang et al. (2024); Mitra et al. (2024); Mondal et al. (2024); Hu et al. (2024a).
@@ 2 Related Work
CoT has been adapted to the vision and language setting Zhang et al. (2024) by adding external knowledge Mondal et al. (2024), extracting a scene graph Mitra et al. (2024), or using visual sketches as intermediate reasoning steps Hu et al. (2024a). Most existing works focus on benchmarks such as ScienceQA Lu et al. (2022) and visual commonsense reasoning Schwenk et al. (2022), with (a) definitive/objective answers; and (b) simple evaluation metrics (e.g., ScienceQA is multiple-choice). We focus on multimodal explanation generation tasks in which the answers are open-ended and nuanced. As in CoT, we elicit intermediate reasoning steps from the models, but propose a novel method using the information bottleneck principle to guide generation and selection of useful knowledge for a correct explanation.
@@ 3.1 The Information Bottleneck Principle
where $I$ denotes mutual information, and $\alpha$ is a parameter to balance compression term I(S,S^)I($\hat{S}$) with relevance term $I(\hat{S},Y)$.
@@ 3.2 Eliciting Multi-Hop Implications
To guide implication selection for generating the correct output, we provide the image $I$ and caption $C$ to the VLM to generate a set of candidate explanations at each hop: $R^{h}=\{r_{1}^{h},r_{2}^{h},\dots,r_{k}^{h}\}$. One candidate explanation acts as an initial hypothesis, refined iteratively when additional information (implications) becomes available. In the first hop, we generate candidate explanations by providing the VLM with the image $I$, caption $C$, and descriptions $D$. From the second hop onward, we condition—in addition to the previous inputs—on each of the $k$ implications selected in the previous hop (§3.3) to generate $k$ candidate explanations. The prompts used for generating image descriptions, implications, and candidate explanations are in Appendix F.
@@ 3.3 Selecting and Refining Useful Implications
The relevance term is used to ensure that implications provide useful information for generating a target explanation. Since our method is unsupervised, we use the VLM to generate candidate explanations at hop h1h-1: $Y=\{R^{h-1}\}_{i=1}^{I}$, which we use as a proxy for the gold standard answer in the next hop $h$. We measure the relevance of the target implication $P_{j}^{h}$ as the maximum probability (minimum cross entropy loss) for predicting the candidate explanation from the current (textual) inputs $\hat{Z}_{j}^{h}=\{C,D,P^{h-1},P_{j}^{h}\}$, which include the caption, image description, implications from previous hops, and the target implication:
@@
where $\alpha$ is a hyperparameter that controls the trade-off between the compression and relevance terms. In our experiments, we set $\alpha=0.7$, based on our empirical observation. A detailed analysis of the effect of varying $\alpha$ is provided in Appendix E.
@@ 3.4 Generating Final Answer
We used Sentence Transformer [^2] BAAI/bge-large-en-v1.5 for all sentence embeddings. The prompts for generating multi-hop implications and explanations are in Appendix F.
@@ 4.3 Baselines
ZS generates a final explanation directly from the image and caption using VLM. CoT follows a similar setup but instructs the model to produce intermediate reasoning chains Wei et al. (2022). Additionally, we implement SR, a multimodal variant of self-refinement Madaan et al. (2023), where a generator produces a response, and a critic evaluates it based on predefined criteria. The critic’s feedback helps refine the output iteratively [^4] Refinement steps set to 2 for fair comparison.. Evaluation criteria include correctness, soundness, completeness, faithfulness, and clarity (details in Appendix H). SR-noC functions identically to SR but without a critic model, refining candidate explanations without feedback. This also serves as an ablation of the implications from BottleHumor. Prompts for baselines are in Appendix H.
@@ 4.4 Evaluation Metrics
For recall, we decompose the reference refref into atomic facts: $\{y_{1},y_{2},...,y_{n}\}$ and check whether each appears in the predicted response predpred.
@@
where $n$ is the number of atomic facts in refref.
Precision follows the same process in reverse, decomposing predpred into a list of atomic facts: $\{x_{1},x_{2},...,x_{m}\}$ and verifying their presence in refref:
@@
where $m$ is the number of atomic facts in predpred. Both decomposition and verification use Gemini-Flash-1.5 with a temperature of 0.2.
@@
To assess the reliability of our metrics, we conducted a human evaluation on 130 random samples across all models and datasets via CloudResearch (details in Appendix D). Human annotators determined whether each atomic sentence appeared in the corresponding text (e.g., reference). The average agreement between the LLM-based evaluator and two human annotators was 77.1% ($\kappa=54.1$), similar to the agreement between the two annotators: 75.4% ($\kappa=50.8$), indicating considerable alignment with human judgment. Prompts are in Appendix G.
@@ 5.2 Contribution of Individual Components
Table 2 presents an ablation study where only a single input is provided after refining implications and candidate explanations. GPT-4o and Phi perform better with both inputs, suggesting they effectively integrate relevant information from both to generate improved explanations. In contrast, Flash-1.5 and Qwen2 models rely more on the candidate explanations, which contain more readily-useful information than the implications, indicating these models are less proficient at ignoring noisy or irrelevant implications.
@@ 5.3 Assessment of the IB Framework
Table 3: Precision, Recall, and $F_{1}$ scores on intermediate explanations across hops. h stands for hop. In our experiments, hop $h=0$ corresponds to $k=0$ (no implications), $h=1$ allows up to $k=3$ implications, and $h=2$ allows up to $k=6$ implications.
@@ 6 Conclusions
Trade-off between interpretability and efficiency.
Our method emphasizes interpretable, step-by-step controllable reasoning for the humor explanation tasks, but this comes with increased resource cost. While the computational cost can be managed by limiting the number of implications or image descriptions, the increased cost remains an inherent trade-off for incorporating interpretable reasoning steps. In contrast, less interpretable or controllable approaches may offer greater efficiency. Each call typically involves $\leq$500 input tokens and $\leq$128 output tokens, with up to 20 calls per sample. For 100 samples, this results in an estimated total cost of up to \$4–5 USD using GPT-4o and up to \$1 USD using Gemini-Flash-1.5-8B.
@@ Appendix D Details on human anntations
Appendix E Effect of $\alpha$ on explanation quality.
To understand the role of the hyperparameter $\alpha$ in Eq. (1), we conducted an ablation study evaluating F1 scores from GPT-4o across all three datasets (100 samples each), using a fixed random seed. As shown in Table 4, performance varies with different $\alpha$ values. When $\alpha=0.0$, the model prioritizes relevance alone; when $\alpha=1.0$, it focuses exclusively on compression. Across datasets, the best performance tends to occur near $\alpha=0.7$, indicating that a balanced trade-off between compression and relevance yields the most informative and accurate intermediate explanations.
Dataset $\boldsymbol{0.0}$ $\boldsymbol{0.3}$ $\boldsymbol{0.7}$ $\boldsymbol{1.0}$
Memecap 46.7 51.5 51.5 48.2
NewYorker 57.5 57.6 57.7 57.1
YesBut 58.1 56.2 59.0 55.0
Table 4: Ablation study for the hyperparameter $\alpha$ used in Eq. (1).
Appendix F Generation Prompts for Selection and Refinement
@@
Appendix G Evaluation Prompts
@@
Appendix H Prompts for Baselines